FEL Upgrade Project Weekly Brief
February 10-14, 2003
 
 
Highlights:
This week was an important milestone for FEL commissioning activities. We had an important institutional goal of completion of the qualification and installation activities for the first light configuration this week because we lose our access to installation, alignment and vacuum labor which is needed for the CEBAF shutdown for the next 6 weeks. With the installation today of the last large complicated vacuum chamber (the fourth optical chicane chamber) and with the second optical cavity assembly going under vacuum we met this essential goal.

Commissioning Activities:
Beam transport:
The fourth and last optical chicane chamber was vacuum qualified this week after several days of a hot nitrogen purge and was installed today. Only one simple field fit vacuum pipe needs to be installed to connect the wiggler output to the outcouper optical cavity and the first light configuration is in place This pipe was left off until last to allow access to the area within the ring. It will be installed early next week. The last two magnets (2 GQs) needed to complete the second arc and the high power light (recirculation) configuration arrived this week. All components (5 dipoles and their chambers) are now on site for the second arc and will be installed late in March.
Optics:
The high reflector optical cavity (the outcoupler) was completed and it goes under vacuum by this weekend with the start of a hot nitrogen purge.
Photoelectron Gun:
A parallel process was begun this week to get the gun back on line as soon as possible and analyze and repair the failures we have noted in our field emission coatings on the electrode structure. In parallel with the electrode repairs, the gun ceramic assembly will be put back under vacuum today so that this entire assembly and the gun HVPS can be tested next week to full operating voltage (500kV). (see more details in Injector section below)

Special Meeting Notices:
FEL Users and Laser Processing Consortium Workshop: March 19-20, 2003
Our annual users meeting will be held on March 19-20 at the CEBAF Center Auditorium at Jefferson Lab. This meeting is open to anyone interested in past and planned future use of the FEL User Facility for basic, applied and mission related R&D. The preliminary agenda for this meeting will be posted on the JLab WEB site next week.

BESAC Facilities Review: February 21-22, 2003
Jefferson lab has been invited to make a presentation at a subcommittee meeting of the DOEs Basic Energy Advisory Committee (BESAC) that will be reviewing potential new facilities or facility upgrades of interest to the BES community. We will report on potential basic science applications of the light source technology demonstrated by the 1 kW and 10 kW FEL devices. (Meeting open to DOE BES invitees only)

FEL Upgrade Cost and Schedule Review: March 24, 2003
At Jefferson Lab (ONR invitees only)

FEL IR/UV Upgrade Semiannual Review: April 9-10, 2003
At Jefferson Lab (ONR/AFRL/DOE invitees only)

Management:
Preparing for the above noted meetings and carefully programming the transition from the installation activities that were completed this week.

WBS 4 (Injector):
The gun team met several times this week to continue several parallel paths to get the gun back in operation as soon as possible, analyze the cause of coating failure observed on the electrodes after the n240 kV tests, and prepare new electrodes.

First of all a new mating flange was prepared for the gun ceramic stack assembly so that the ceramic stack could go under vacuum today. Next week we plan to a full voltage test (up to 500 kV) of the gun HVPS and ceramic stack assembly.

Secondly, the electrode assembly that was conditioned up to 240 kV will have its failed coating removed. Any areas that need touch up polishing will be polished to our standard 1 micron finish and we will re-install this electrode assembly back into the gun in order to get the gun/injector back up in March for first light tests.

Simultaneously, we can fix the problems we identified in the first gun conditioning runs (misaligned seating of the GaAs crystal which caused it to crack, the door mechanism, and flaky ion pump and BPMs in the injector line). We do not feel that there is any significant risk in bringing the gun back on line with the uncoated electrode assembly if we limit the voltage to

350 KV which is all we need for the 10 kW IR Upgrade, because of our (now tested) internal cesiation scheme and the tested charge bleeding resistivity of the implanted gun ceramics.

Thirdly, we are looking at all the process parameters that might have led to the observed coating failure. We believe that the observed coating failure on the cathode ball ( a small ring of etch pits around the GaAs location) is probably due to the excessive field emission due to the original cracked crystal. The observed (at 10x) failure of the coating near the low field end of the cathode support tube points to a fundamental adhesion problem either due to thermal stress from the bakeout or an insufficient pre cleaning of the substrate by the pre-deposition plasma cleaning step. A number of new coating samples will be produced using our best known process parameters and these samples will subject to adhesion and thermal stress tests. When a benchmarked coating process is re-obtained we will coat our back-up set of electrodes (from the 1 kW gun) and re-install these electrodes later in the year. Performance of the gun at higher voltages (>350 kV) is not needed until we plan to qualify the gun for UV FEL operation in the summer. We thank the extra efforts that are being provided by the William and Mary thin film team under the direction of Dennis Manos to help us solve this problem.

The electrode structure was disassembled and inspected to allow a determination of what caused the low voltage field emission in the gun assembly last week. We found damage to the field emission coating near the cathode that may have been caused by field emission from the cracked cathode earlier. The damage is concentrated close to the cathode position and not the high field point of the cathode ball. A few arc spots were also found at the high field point of the cathode support tube and the coating on the support tube was bubbling and delaminated in spots. Silica dust from the exfoliated coating was found near the cathode ball. The group of people who coated the electrodes has been assembled to study the damage and determine the cause of the coating damage. We are working on setting up to recoat the electrodes using a recipe that previously demonstrated the best results. A test run with a test tube will be done first to qualify the recipe. In parallel with work on the electrode, the leak in the stack was isolated to a flange seal and repaired. The stack will now be assembled without the electrode and run up the full voltage of the high voltage power supply to qualify the ceramic stack and the high voltage power supply. We will also be testing previously coated electrodes to eliminate some other possible causes of coating failure.

WBS 6 (RF):
The RF is operational throughout the machine.

Injector - Operated and tuned both Quarter cavities and the Buncher.

Zone 2 - Corrected a download file to allow the correct GMES reading. A faulty Arc Detector test was traced to a loose cable and corrected. The zone was operated at the Drive High gradients.

Zone 4 - A faulty crowbar was repaired by replacing a power diode, an SCR, and a 24 volt regulator IC. SRF is setup to check the cavity calibration values. This is now scheduled for next Tuesday and Wednesday.

WBS 8 (Instrumentation):
The Machine Protection System (MPS) commissioning is progressing well. All expert screens are operational and the high level operations screens are nearing completion. The RF systems "P1" faults are connected and tested. These and the Beam Loss Monitor (BLM) outputs are the "fast" inputs to the MPS, they are connected to the system via fiber optic cable with the permit signal being a 5 MHz pulse train. All four of the BLM boards have been powered up and are communicating with EPICS. The basic functionality is being verified and the higher level functions such as trip sequence logging are being developed. Also, the high level EPICS screens are being updated. A current snap-shot of the BLM-overview screen and a photo of the four boards running in iocfelt2 in user lab 5 are included as this weeks pictures at http://laser.jlab.org/wbs8. For those who are interested, a link to the BLM design review is included there too. The design for the magnet current and switch position MPS interface chassis is now complete. Two chassis are being fabricated while the PCB layout is in progress. These accept up to six analog inputs (magnet currents) and 8 digital inputs (magnet switch positions) and convert the signals to 24 volt levels for determining valid locations for the electron beam (Machine Modes of the MPS).

Video System: Minor maintenance and quick tasks for the video system are always a standard part of each week. In addition to fixing cabling issues for user lab 6, the online diagnostic video channel has been has been changed in the video switcher for temporary remote monitoring of the optics work in the vault over the week-end. The link is: http://laser.jlab.org/video.

Core winding termination for the sextupoles in the 1st Arc is complete. Shake out and commissioning of the Dispersion and Wiggler magnets was successfully completed this week. Routing of DC power for the extraction GV magnets and trims has begun. The preliminary design for the QT/GC vertical correctors is ready. We will use Apex PA12A power OpAmps combined with the new "Analog Block" being developed by the EES group to create the 8 channels of floating 5Volt / 2 Amps.

Installation of the 2G00 and 2G00A beamviewers was completed. Additional Beam Viewer Assemblies have been assembled and camera assembly is in progress. New patch panels are being made for video and power to experiments and equipment in the vault. Work continues on the paddle viewer foils, this time Dave Waldman is giving it a try.

The FEL Safety Assessment is in progress. Thanks to all who have donated their time to help this effort. Drawings were completed for the new Dump Current Buffer Driver (Version 2) schematic. Parts have been ordered and assembly will begin as soon as the new PCBs are received.

WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in the First Arc.
The thermal protection circuitry was wired in anticipation of an operational test.
Dipoles
Optical Chicane Dipoles (GW)
No progress.
Injector Dipoles (DU/DV)
Jeff Dale has the second GU apart in a process of re-gluing the Purcell gap plates.
Arc 180 Degree Dipoles (GY)
In preparing the GY for measurement, we discovered that the glue joint in the isolated end Purcell gap shims in the GG coil and also main body are peeling. Furthermore we found up to 0.010 inch rises in the steel shim at some of its joints. We have determined an approach to fixing these problems in adherence and have initiated the repair activity.

In the mean time, all other modifications and preparations on the magnet are complete. Left side field clamps are expanded, field clamp adjustment modifications done, hoses are complete, the manifold bracket of the GG coils changed for better hose placement and the steel bulkhead fittings in the GG coils changed to brass and fitted to the leads of the GG coil.
Wang potted the second coil of the second GY.
Arc Bend & Reverse Bend Dipoles (GQ, GX)
We received the remaining two GQ magnets on Tuesday to complete the Bend-Reverse Bend (GX-GQ) order.
Wang NMR is making the replacement GX coil for the #2 GX. Magnet.
Quadrupoles
Trim Quad (QT)
The two QT quads with the coils that create the GC vertical corrector were tested for corrector field goodness. The part in 10 field integrals from the corrector doesnt meet Daves specification of a part in 100 so they are rejected. However, David Douglas allowed these vertical correctors for the arcs to be released to the enclosure (as we did in the IR Demo) as an experimental magnet system. Beam distortion from large correction values from these magnets can be avoided after they are reduced in excitation after a realignment of the machine. In the mean time Tom Hiatt is planning to raise current density in his magnetic model of the add on coils as a trial method that may remove some of the field integral droop at the sides of the magnet.
Sextupole (SF)
The Sextupole "lights" were wired up in the first arc.
Octupole (OT)
The order is ready to be placed anticipating the next batch of commissioning money.
Beam Line and Vacuum
The vacuum group released the last chamber in Optical Chicane after they baked and purged to remove unexpected oil contamination. The chamber was installed in the magnet this morning.
The thermal protection in magnets of the 1st Arc was wired up in anticipation of a commissioning test.
The 180° Bend chamber for the second arc continues to await the receipt of the diagnostic windows from Ceramaseal. The latest projected delivery date is late February.
The Vacuum Group is working on the temporary vacuum pipe between the temporary dipole after the wiggler and the output can of the optical cavity that will allow first light operation.

WBS 10 (Wiggler):
All connections to the optical klystron were completed and the control software checked out. The plexiglass cover required by OSHA was installed. There is still a short punch list and the shunts still have to be checked out but the optical klystron is now ready for beam.

WBS 11 (Optics):
On the HR, the UHV water connections were completed and the 220 L/s ion pump installed. Leak checking found one leak at a gasket, this was replaced. No other leaks were found. A hot nitrogen purge is running over the weekend to remove water from the insulation. Early next week we will install the tube that connects the optical beamline to the electron beamline. On the OC, several wiring issues are being resolved. Today we plan to complete the water cooling loop, and do vacuum work on the diagnostic hardware just upstream of the vacuum vessel.

Other activities:
The LSOP for User Lab 6 was modified to include the laser hardware purchased by several of our users. We resolved some ambiguity in the current-carrying capacity of the 3-phase power, and the wiring and connectors are being changed.

An oral presentation on our proposal to produce laser damage resistant optics or high power FELs was prepared to send to the JTO.